if y^2+16y-4x+4=0 then what is the equation of the directrix?
Answer by lwsshak3(11628)

if y^2+16y-4x+4=0 then what is the equation of the directrix?
..
y^2+16y-4x+4=0
complete the square
y^2+16y+16=4x-4-16=4x-20
Standard form: (y-k)^2=4p(x-h)
(y+4)^2=4(x-5)
This is a parabola with vertex at(5,-4), and axis of symmetry, y=-4
4p=4
p=1
ans: equation of directrix, x=4
see graph below
..